OpenStudy (anonymous):

Use the distributive property to simplify the expression below: 5t − 7(3 + t) a. 12t-21 b. 12t-14 c. -2t-21 d. -2t+21

OpenStudy (zehanz):

Distributive property: a(b + c) = ab + ac. Why not use it?

OpenStudy (anonymous):

so the a=5t-7 and the 3+t is the b+c?

OpenStudy (zehanz):

No, leave out 5t for now, we'll deal with that later. a = -7, b=3 and c = t

OpenStudy (anonymous):

okay so -7 X 3 + -7t ?

OpenStudy (anonymous):

-21 + -7t..?

OpenStudy (zehanz):

Yes, and -7*3=-2. Now, remember you also have 5t lying around somewhere...

OpenStudy (anonymous):

how is -7*3 = -2? shouldn't it be -21?

OpenStudy (anonymous):

...

OpenStudy (zehanz):

Of course, it is a typo, sorry!

OpenStudy (anonymous):

lol okay, so now i have -21 + -7t right? and a 5t laying around somewhere..?

OpenStudy (zehanz):

Right: 5t -7(3+t) = 5t-21-7t=...

OpenStudy (anonymous):

-2t-21 so.. c

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (zehanz):

YW!